OpenStudy (anonymous):

A circular jogging track is being designed so that one complete lap around the track is one mile long. What should the radius of the track be to the nearest foot?

OpenStudy (anonymous):

HINT... perimeter of the circle is... \[\LARGE P=2\pi r\] you have perimeter 1 mile... you have pi =3.14 find radius

OpenStudy (anonymous):

UHHH

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

yup im doing it but me too lazy lol

OpenStudy (anonymous):

haha.... just try :P

OpenStudy (anonymous):

GD i just broke my pencil :P brb lol

OpenStudy (anonymous):

\[\LARGE \frac{P}{2\pi}=r\] hahahah ... lol now substitute :P

OpenStudy (anonymous):

can i haz answer :P lol

OpenStudy (anonymous):

ehhh... try it . it's easy !

OpenStudy (anonymous):

.

OpenStudy (anonymous):

841 FEET